Fact Pattern:

National Industries is a diversified corporation with separate and distinct operating divisions. Each division’s performance is evaluated on the basis of total dollar profits and return on divisional investment.

The WindAir Division manufactures and sells air-conditioning units. The coming year’s budgeted income statement, based upon a sales volume of 15,000 units, appears below.

WindAir Division

Budgeted Income Statement

For the Next Fiscal Year

Per

Total

Unit

(000 omitted)

Sales revenue

$400

$6,000

Manufacturing costs

Compressor

$  70

$1,050

Other materials

37

555

Direct labor

30

450

Variable overhead

45

675

Fixed overhead

32

480

Total manufacturing costs

$214

$3,210

Gross margin

$186

$2,790

Operating expenses

Variable selling

$  18

$   270

Fixed selling

19

285

Fixed administrative

38

570

Total operating expenses

$  75

$1,125

Net income before taxes

$111

$1,665

WindAir’s division manager believes sales can be increased if the unit selling price of the air conditioners is reduced. A market research study conducted by an independent firm indicates that a 5% reduction in the selling price ($20) will increase sales volume by 16%, or 2,400 units. WindAir has sufficient production capacity to manage this increased volume with no increase in fixed costs.

WindAir currently uses a compressor in its units that it purchases from an outside supplier at a cost of $70 per compressor. The division manager of WindAir has approached the manager of the Compressor Division regarding the sale of a compressor unit to WindAir. The Compressor Division currently manufactures and sells a unit exclusively to outside firms that is similar to the unit used by WindAir. The specifications of the WindAir compressor are slightly different, which will reduce the Compressor Division’s materials costs by $1.50 per unit. In addition, the Compressor Division will not incur any variable selling costs for the units sold to WindAir. The manager of WindAir wants all of the compressors it uses to come from one supplier and has offered to pay $50 for each compressor unit.

The Compressor Division has the capacity to produce 75,000 units. The coming year’s budgeted income statement for the Compressor Division is shown below and is based upon a sales volume of 64,000 units without considering WindAir’s proposal.

Compressor Division

Budgeted Income Statement

For the Next Fiscal Year

Per

Total

Unit

(000 omitted)

Sales revenue

$100

$6,400

Manufacturing costs

Materials

$  12

$   768

Direct labor

8

512

Variable overhead

10

640

Fixed overhead

11

704

Total manufacturing costs

$  41

$2,624

Gross margin

$  59

$3,776

Operating expenses

Variable selling

$    6

$   384

Fixed selling

4

256

Fixed administrative

7

448

Total operating expenses

$  17

$1,088

Net income before taxes

$  42

$2,688

Question

What is the effect on the Compressor Division’s net income of the sale to WindAir of 17,400 compressors if some external sales must be lost?
Answer #1

CALCULATION OF VARIABLE COST

SITUATION I II
SELL OUTSIDE FIRM SELL TO WIND AIR
RAW MATERIAL 12 12-1.5=10.5
DIRECT LABOUR 8 8
VARIABLE OVERHEAD. 10 10
VARIABLE SELLING 6 0
36 28.5

CALCULATION OF CONTRIBUTION

I II
SELLING PRICE 100 50
(-) VARIABLE COST 36 28.5
CONTRIBUTION 64 21.5

FIXED COST IN BOTH THE SITUATION WILL REMAIN SAME AND HENCE NOT RELEVANT

TOTAL CONTRIBUTION OF COMPRESSOR NOT SOLD TO WINDAIR:

UNITS* CONTRIBUTION=64000(OUTSIDE)*64= 40,96,000..............(1)

TOTAL CONTRIBUTION IF 17400 COMPRESSOR SOLD TO WINDAIR

UNIT 17400(WINDAIR) 75000-17400=57600(OUTSIDE)

* CONTRIBUTION * 21.5 (AS CALC. ABOVE) * 64

374,100 + 36.86,400

Total= 40,60,500.................(2)

comparing (1) and (2) , we get (35,500)

Hence there would be an decrease in net income 35,500